OpenStudy (anonymous):

input output Which rule describes how the 0 -3 output depends on X? 1 -1 2 5 A) 2x^2-3 3 15 B) -3x+2 C) 2x-3

OpenStudy (anonymous):

help me someone pleaaaase.

OpenStudy (ivettef365):

Substitute the x for the value of x do the problem to see if the result is the value of y

OpenStudy (ivettef365):

For example you have 2x^2 -3, when you plug 0 it works because the result is -3,when you plug 1 for x then you ha 2-3 = -1, it works, when you plug 2 then you have 8-3=5

OpenStudy (anonymous):

thank you

OpenStudy (ivettef365):

what about when you use 2x-3, with 0 it works because you have 0-3= -3, when you plug 1, then you have 2-3=-1, but when you plug 2, then you have 4-3= 1, then does not work